A refrigerator contains 5 mangoes, 4 oranges and 3 bananas.

a. if you pull out a piece of a fruit art random, what is the probability you get a mango?

b. what is the probability that you do not get a mango?
Answer #1

There are total 5 + 4 + 3 = 12 fruits.

a)

P(mango) = n(mango) / Total fruits

= 5 / 12 ( = 0.4167)

b)

P(Not mango) = 1 - P( mango)

= 1 - 5/12

= 7/12 ( = 0.5833)